OpenStudy (anonymous):

a student guesses the answer to 6 question on a true-false quiz find the probability that the indicated number of guesses are correct :exacly 4

jimthompson5910 (jim_thompson5910):

Use the binomial probability distribution formula P(X = x) = (n C x)*(p)^(x)*(1-p)^(n-x) P(X = 4) = (6 C 4)*(0.5)^(4)*(1-0.5)^(6-4) P(X = 4) = (6 C 4)*(0.5)^(4)*(0.5)^(6-4) P(X = 4) = (15)*(0.5)^(4)*(0.5)^2 P(X = 4) = (15)*(0.0625)*(0.25) P(X = 4) = 0.234375 So the probability of getting exactly 4 questions correct is 0.234375

OpenStudy (anonymous):

thank you so much.

jimthompson5910 (jim_thompson5910):

you're welcome

OpenStudy (anonymous):

give the value of the expression 11C4 a)330 b)220 c)110 d)550

OpenStudy (anonymous):

give the value of the expression 11C4 a)330 b)220 c)110 d)550

jimthompson5910 (jim_thompson5910):

n C r = (n!)/(r!(n-r)!) 11 C 4 = (11!)/(4!*(11-4)!) 11 C 4 = (11!)/(4!*7!) 11 C 4 = (11*10*9*8*7!)/(4!*7!) 11 C 4 = (11*10*9*8)/(4!) 11 C 4 = (11*10*9*8)/(4*3*2*1) 11 C 4 = (7920)/(24) 11 C 4 = 330

OpenStudy (anonymous):

log of base 9 +log base of 2 a=log base 2 11

jimthompson5910 (jim_thompson5910):

Can you draw this one out?